OpenStudy (anonymous):

The mass of Jupiter is 1.908 x 10^24 metric tons. The earth has a mass 6 x 10^21. How many times bigger is Jupiter?

OpenStudy (anonymous):

198.46 times. Take (1.908*10^24)/(6*10^21)=198.46

OpenStudy (anonymous):

If you need scientific notation, it's 1.9846*10^2

OpenStudy (anonymous):

Thank you, however how did you get 1.9846 I thought you multiply 6 x 1.908 to get 11.448

OpenStudy (anonymous):

Division is a ratio. Divide the mass of Jipiter by the mass of Earth.

OpenStudy (anonymous):

so it would be (1.908 x10^24) / (6 x 10^21)

OpenStudy (anonymous):

Exactly

OpenStudy (anonymous):

ok, but I get .318 x 10^-3

OpenStudy (anonymous):

Check your math. The rule of exponents will allow us to subtract 10^24-10^21.

OpenStudy (anonymous):

ok then I have .318 x 10^3 right

OpenStudy (anonymous):

Nope, that would be 318.

OpenStudy (anonymous):

Let's try this...

OpenStudy (anonymous):

(1.908/6)=.198

OpenStudy (anonymous):

.198*10^3=198.46

OpenStudy (anonymous):

Before subtracting the exponents, simplify the rest of the fraction first.

OpenStudy (anonymous):

I understand the divide the mass of Jupiter by the mass of Earth. (1.908/6) but I get 0.318. I just don't get it. I ams sorry. But thank you very much for trying to explain it to me.